Search

Reports courtesy of Saturday (Mike Hallam-Rudd) and Sunday (Beth Marsh):


Saturday saw an unusually low number of volunteers with only 5 present. Peter Vickers continued his systematic project on checking out wiring circuits on 15, this is one of those tasks that can be a quick fix if all the systems bell out OK or Not. in the latter in can then lead to lots of checking. Today he was working on the conduit run for the No 2 Cooler Group Fire Warning system, No 1 end still to do. Graham was continuing working of wiring removal from underneath No 9. He is slowly but surely getting there, I think today was one of those 'slowly' days judging by the comments coming out of the pit.

Geoff starting the day doing some touch up painting in No 1 end of 19 before going to the No 2 end to add a few touches there. Bob is picking up on this tomorrow and working through the air tanks after they have been installed. Mike and Richie started the day recovering a pipe union at No 2 end which had slipped down the pipe and gone through the pipe hole and under the cab floor, result👍, cab floor up, 20 seconds later Richie was putting it back down again. Richie secured the tank attached to the drivers seat pedestal whilst Mike went to prep tools and the hoist at No 1 end.

Their attention then turned to No 1 end firstly with the installation of the 2 x Co2 Fire Bottles. Richie then secured the tank which goes under the Traction Motor Blower after Mike had connected one of the end pipes. The Vacuum Exhauster table was next, these have been known to take a long time, on one occasion over a day, trying to line up the 'sprung' legs using the adjusters. We managed it in under 1 hour with no lost tools or finger ends, the plumbing on this tank was then completed. Geoff then assisted Mike after Richie had departed in going through the days work ensuring that all fittings we had worked on were suitable tight, it would be difficult to access some parts once all is built up. We then tidied the clutter in both cabs.


Turning to Sunday:

55019 - Work progressed with rebuilding No.1 nose end, the traction motor blower table was refitted and a small air tank just in front of that was also refitted. This was followed up by a paint, touching up the air tanks and exhauster table that were previously refitted prior to today.

D9015 - Work continued with replacing the flexible pipework in the radiator bays, 3 bays are now fully complete with 5 more to go. A couple of bays were completed last week but awaiting some different diameter pipe to finish off.

No.8's Cab - We started off with undercoating the roof section, using the tower scaffold we got into place last week. It was then time for Blue, from the yellow nose down to the cab doors both sides was given a top coat of BR Blue. The cab side doors were also top coated in blue as were the nose end hatch doors. Work also continued with internal jobs getting the parts for the doors ready to refit. The side door roller tracks were degreased cleaned up and some primer applied on the mounting squares to prevent more rusting. The train heating box which sits in the middle of the windows was given some attention over the past few weeks, The outside was stripped down and primed previously and today the same treatment was given to the inside.



Updated: Jul 12

Mike Hallam-Rudd, DPS Volunteer Liaison Officer, provides a report from the Saturday team covering May & June 2021:

First of all apologies for the lack of this updater for the last couple of months. Although this report covers May and June and is just an overview. Although the air tanks were removed in late April some of the background work for the preparation was carried out earlier in April and also thanks to the support of the Sunday crew who have done some of the mundane and unseen tasks such as cleaning and painting inside the nose ends of 55019 and the equipment which was removed from them to enable the air tanks to be removed and sent away to Tecforce at Derby for their 10 Year Certification test. 1 of which failed with the threads on one of the boss holes being below spec but a repair was authorised and the tank will be returned back to the Depot on Friday 9th July. Fortunately it was 1 of some 4 similar tanks so we can fit this one in last in No 1 end above the TMB.

Much of the work carried out during this period had been on 19 to get it ready for its visit to the Great Central Railway in August. It will stay there until 9 arrives and the GCR are kindly allowing us to do a Double Deltic running weekend, please watch both parties’ websites and FB pages for updates, this is likely to happen in October. The loco will arrive a couple of days before and the trailer will remain at Quorn and return 19 to BH on the Monday.

The Power Unit team, lead by Pete Stojanovic have given some valuable help to the task on 19 due to the move date being bought forward 2 weeks due to a joint move deal with the Heavy Tractor Group and their 37714 returning to the GCR on the 2nd August and Allelys Heavy Haulage who have agreed to a good commercial deal to then collect 19 the following day with the same rig. Because of Covid restrictions and workspace inside the nose ends some of the PU team were able to continue with work on unit 417 which is destined for 9 and replacing 538, ex D9015. There is about 2 days work left to complete this task.

During May we had a visit from Steve Bradley of Merritts Machine Movers, Nottinghamshire who is doing the PU swap for us and he has also quoted to do the lift on 15 in connection with the bogies should it be needed as we currently have no access to a set of Jacks.

Graham has been continuing to work on removing the defective cables from 9 and we have now sourced a contractor, member and volunteer Steve Nutter has his own company Envirolectrics and he is releasing one of his staff to assist for 1 day per week, and hopefully 2 days if he can. Another positive in 9’s path to return to traffic was Saturday 26th June when Graham found the test schedules which our other contractor needed to test some of the other items delivered back on April 28th (well it’s nearly in May). This will massively help speed up this task. A couple of volunteers have also offered to assist Graham midweek where required.

Meanwhile, still on electrics Peter Vickers has been working away doing circuit testing on D9015 and is slowly but surely getting there, most good news but he has identified a few things which need to be replaced. Must not forget Geoff who has been busy working on some electrical stuff at home.

The last couple of weeks has seen the start of the rebuild of 19’s air system. We also had the pleasure on the 26th June of hosting a Armed Forces Day mini event amongst the team courtesy of Richie Conley. The 2 tanks under the battery boxes were completed today 3rd July by Pete with help from son Alex who also cleaned off the dirt from the battery tops and terminals. Mike and Fiona completed the fitting out of No2 nose end. A team is going down on Tuesday 6th July to make a start on No 1 end, the small tanks attached to the cab seat pedestals will also be done and any snagging work and outstanding tasks will be completed on the 10th.





Subject to no further delays in the lifting of Covid-19 restrictions, it is planned that Barrow Hill will be reopening to the public on 7th and 8th August.


From the reopening, the DPS depot will now be open on Saturdays and Sundays.

To receive e-mail notifications each time a new post is made here, please click the Login/Sign up button to the right. Once you have an account, visit the Settings area to subscribe or unsubscribe from Blog subscriptions.